

********************************************************************
UARS WEEKLY STATUS REPORT - 17 SEP 2004
********************************************************************
**********************************************************
EMAIL from UARS Project Science Office (15 September 2004)
**********************************************************
Dear UARS Investigators:
Thirteen years ago yesterday, September 14th, the UARS was deployed from
the Space Shuttle Discovery. We consider it our lucky "13th Anniversary
of Deployment."
The EOS Aura satellite was launched successfully in July of this year
and we look forward to UARS providing at least a year of overlap of
measurements to help in the validation and calibration of the new
EOS Aura atmospheric-observing instruments.
Five of the original ten UARS instruments collect data on a regular basis.
These five instruments include HALOE, SUSIM, SOLSTICE, HRDI, and PEM.
Although there have been some recent battery concerns, the skillful and
professional UARS Flight Operations Team is in control of the situation
and adjusts the battery charging as needed.
We salute the many personnel who have made the UARS mission possible:
Those who designed and built a spacecraft and instruments capable
of lasting 13 years in space.
Those who monitor the functionality of the instruments and
provide daily activity plans.
Those who monitor the spacecraft 24/7 and have overcome several
very serious anomalies over the years.
and Those who have harvested the 13 years of data providing
breakthroughs in our scientific understanding of many processes
related to the atmosphere.
Thanks to all for your support over this many-year mission.
Best regards,
Charles Jackman & Anne Douglass
UARS Project Science Office
********************************************************************
***********
SOLSTICE
***********
This week, SOLSTICE continued to operate in good health gathering solar
data on all available orbital tracking periods. On average, we received
55% of each day's data amounting to about 524 minutes of solar
observing time per day.
***********
SUSIM
***********
As available telemetry time permitted, SUSIM successfully gathered
daily mid (1.1 nm) resolution, MgII, occultation, solar pointing offset,
and continuous monitoring scans. SUSIM operations were unusually and
severely limited on several days during this past week because of the
lack of telemetry time.
***********
PEM
***********
PEM wishes to congratulate everyone past and present for the 13 years
of successful operations of the UARS!
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004099 has been reprocessed and archived at the GSFC DAAC.
MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 2004099.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing recieved.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 04 September 2004 (Orbit 71044, GMT Day 248)
through 10 September 2004 (Orbit 71149, GMT Day 254).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 250/1401 71083 POWER ON
HALOE 251/0641 71094 POWER OFF
HALOE 251/1539 71099 POWER ON
HALOE 252/0420 71107 POWER OFF
The beta angle ranged between the angles of 62.9 degrees to 58.1
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 25 September 2004 (DOY 269).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.26 to 1.27 degrees C. Additional
battery performance data for each battery on 03 September and 10
September 2004 are provided in Appendix D.
The clock error ranged from 10.5 to -10.3 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 03
HALOE = 04
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 00
Zero (0) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
On 10 August 2004, during the 223/0121 Z event, VT-6 was commanded due
to the differential voltage being greater than 180 mV for three
consecutive orbits. The differential voltage returned to nominal
values after two orbits subsequent to the change. The spacecraft
remained in VT-6 until approximately 245/1600, at which time it was
commanded back to VT-5. All trending of the power subsystem during
this time looks nominal and the FOT will continue to use VT-6 should
the differential climb again.
A new release of the CMS code to allow for SMA events was delivered to
the backup CMS last week. This new release was built to fix the file
transfer problem encountered in the previous build. The preliminary
testing showed the file transfer issue to be resolved, but the build
still requires all other process testing. The goal for completion is
the end of September.
MEETINGS
The FOT attended the UARS Monthly Status Review on 07 September 2004
to review the significant events for the month of August.
The FOT also attended the HRDI attitude meeting on 08 September2004.
At this meeting it was decided that the CDHF functionality of
generating the FDF attitude file, would be reexamined in an attempt to
reprocess old data on the mini-CDHF. This is being done in the hope
that it will give the HRDI team more accurate attitude knowledge to
possibly correct their attitude discrepancy.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 25 September 2004 (DOY 269).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 10 September 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
71049 248/0650 04-251 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
71083 250/1401 04-252 HALOE POWER ON
71094 251/0641 04-253 HALOE POWER OFF (HATBAT HI)
71099 251/1539 04-254 HALOE POWER ON
71102 251/1918 04-255 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
71107 252/0420 04-256 HALOE POWER OFF (TBAT)
71123 253/0551 04-257 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 10 September 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 10 September 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
71049 248/0655 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
71102 251/1918 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
71123 253/0554 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 10 September 2004
03 September 2004 - GMT Day 247 Beta = 58.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.62 +00.0/+00.0 07.0 0.00 0.0
2 +22.4/-044.8 +3.87/+2.29 +20.8/-13.6 26.4 1.04 16.3
3 +11.2/-044.8 +3.23/+1.66 +20.8/-14.4 26.4 1.03 16.3
10 September 2004 - GMT Day 254 Beta = 58.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.31 +00.0/+00.0 07.0 0.00 0.0
2 -11.2/-067.2 +3.87/+2.60 +22.0/-14.4 26.4 1.03 15.9
3 -22.4/-112.0 +2.92/+1.66 +21.6/-14.4 26.2 1.02 15.9
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 10 September 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-059 FOT 71015 246/0136 N/A OBC/CLK adj cmd problem
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.